### v1

# Radon: a Programming Model and Platform for Computing Continuum Systems

Abstract
Emerging compute continuum environments pose new challenges that traditional cloud-centric architectures struggle to address. Latency, bandwidth constraints, and the heterogeneity of edge environments hinder the efficiency of centralized cloud solutions. While major cloud providers extend their platforms to the edge, these approaches often overlook its unique characteristics, limiting its potential. To tackle these challenges, we introduce Radon, a flexible programming model and platform designed for the edge-to-cloud continuum. Radon applications are structured as atoms, isolated stateful entities that communicate through messaging and can be composed into complex systems. The Radon runtime, based on WebAssembly (WASM), enables language- and deployment-independent execution, ensuring portability and adaptability across heterogeneous environments. This decoupling allows developers to focus on application logic while the runtime optimizes for diverse infrastructure conditions. We present a prototype implementation of Radon and evaluate its effectiveness through a distributed key-value store case study. We analyze the implementation in terms of code complexity and performance. Our results demonstrate that Radon facilitates the development and operation of scalable applications across the edge-to-cloud continuum advancing the current state-of-the-art.
Index Terms:
Computing continuum, distributed programming model, middleware, distributed systems.
I Introduction
Cloud-computing consolidated as the de-facto standard for building large-scale distributed applications [1]. In this approach, applications are developed as a composition of loosely-coupled services. In fact, managed environments offer application-agnostic platforms with advanced deployment and execution models, such as serverless functions [2], managed databases [3], or data analytics platforms [4, 5].
However, with the expansion of Internet of Things (IoT), smart cities, and industrial automation, new requirements have emerged that challenge conventional cloud solutions. Traditional cloud-centric architectures face limitations due to latency and bandwidth constraints when transferring data over long distances. The massive amounts of data generated by edge devices can create bottlenecks in cloud-based computing [6], as increased communication distances lead to delays that may violate application requirements.
Research on edge computing aims to address these new scenarios by exploiting computing resources that are close to the sources of the data. Managing systems with edge and cloud components can become complex, so major cloud vendors offer ways to extend their own cloud-native programming primitives to the edge  [^1] https://k3s.io, https://aws.amazon.com/greengrass/ in order to work with the compute continuum. Furthermore, cloud vendors centralize the decision on the programming paradigms and technologies offered. Thus, developers are frequently constrained by specific programming languages, services, and interaction paradigms, limiting their choices, and reducing flexibility and interoperability among different vendors.
To address these challenges, we propose Radon, a flexible programming model and platform to streamline the development of portable distributed applications that covers the edge-to-cloud continuum. In the Radon programming model, applications are decomposed into isolated stateful entities, called atoms. Atoms communicate by exchanging messages and can be configured to realize different software architectures, interactions, and execution paradigms. Atoms can be reused and composed like building blocks to efficiently construct complex applications.
We develop a lightweight, language-independent, and deployment-independent platform that uses WebAssembly (WASM) as enabling technology to implement this model. Developers write atoms in their preferred programming language, compile them to WASM modules, and execute them on an instance of the platform, which we call a runtime. Taking inspiration from operating systems, we define a common runtime interface that atoms use to interact with the implementation aspects that are outside of the domain logic (e.g., storage, communication). As system calls abstract the interaction with the hardware, this interface abstracts key features used to implement distributed applications. This runtime interface decouples the features offered by the runtime from the way they are implemented. Indeed, Radon supports different implementations of the runtime that can be based on diverse technologies. Consequently, it can adapt to the heterogeneous deployment scenarios of edge-to-cloud environments. Different runtime implementations can be tuned to fit the specific scenario (e.g., edge, cloud, serverless, on-premise) without requiring modifications to the code of the atoms. In this view, atoms can focus exclusively on application logic, delegating infrastructural concerns to the runtime. Meanwhile, the runtime can be optimized for heterogeneous deployments and operating conditions without affecting the atoms that run on top of it.
In summary, our work contributes to the research on developing distributed computing continuum applications by:
  1. 1.
    Introducing the Radon programming model and platform, suitable for large-scale and heterogeneous distributed environments.
  2. 2.
    Presenting a prototype implementation of Radon.
  3. 3.
    Providing an initial assessment of the effectiveness and efficiency of Radon through the implementation of a concrete application – a distributed key-value store
The remainder of the paper is structured as follows. Section II outlines our vision for designing and building Radon. Section III describes Radon’s programming and execution model. Section IV showcases the application of Radon in implementing a distributed key-value store, and evaluates its performance. Finally, Section V reviews related work, and Section VI provides a final overview of the paper, highlighting open research directions.
II Vision
The development of Radon starts from the following key observations.
  • While compute continuum environments become more and more relevant, suitable programming and execution models that consider their specificity are still lacking.
  • Developing applications for the compute continuum requires adapting to highly heterogeneous architectures and hardware platforms.
  • WASM can be used as a portable target that can run efficiently on many hardware architectures.
Usually, WASM is used as part of an application and interacts only with the local machine. However, we see potential in using it as a tool for writing distributed applications, particularly in large-scale computing continuum scenarios. In our vision, WASM can solve the portability challenges inherent to heterogeneous applications, while being extended to provide the necessary tools for developing distributed applications. Furthermore, by defining a specification that encodes these extended features as a set of functions, we can decouple their implementation and enable flexible mixing and matching of different implementations based on specific application requirements.
In designing Radon, we took inspiration from operating systems, where programs rely on system calls to interact with their environment while abstracting hardware details. In this context, we recognize that interfaces such as POSIX or the JVM are too broad and hamper portability, as some deployment scenarios may not allow to fully implement them correctly. We intentionally chose to restrict the set of functions made available to WASM modules, excluding the WASI/WASIX specification, which we deem too broad. Instead, we defined our own interface, which is specifically designed to easily build efficient and scalable distributed applications. As a result, the design of Radon was guided by the following principles.
Distributed first. Radon is built for geographically distributed, heterogeneous environments, typical in computing continuum scenarios. It ensures that applications can run across multiple network locations without relying on a centralized infrastructure. To achieve this, applications are composed of independent entities that interact exclusively through message passing. Rather than abstracting away distribution, this design acknowledges it and simplifies its concerns through an interface implemented by a runtime. Therefore, these independent entities can be easily executed, discovered by one another, and can communicate by leveraging the built-in runtime functionalities.
Modularity. Taking inspiration from the development model of service-oriented computing, where applications are built as compositions of services, Radon offers atoms as composable building blocks for distributed applications. In the large, atoms promote modularity as they can be reused across applications. Internally, atoms can rely on the typical layered modularity provided by libraries: developers can implement libraries to be reused within multiple atoms.
Flexibility. Radon promotes flexibility in terms of programming paradigms and languages. It seamlessly supports the development of applications following different architectural styles and execution paradigms. For instance, atoms may be configured to be statically instantiated as named services or dynamically instantiated to respond to function invocations, akin to a serverless environment. Atoms are programmable in any language that supports execution on WASM, such as Rust, JavaScript, Java, C, C++, and Go. This makes the implementation of atoms flexible, as each programming language comes with specific paradigmatic choices.
Portability. The Radon runtime offers a common substrate for atoms, making them portable across highly heterogeneous environments. Indeed, the lightweight nature of the runtime makes it suitable to run on diverse hardware architectures, from resource-constrained ones typically found at the edge of the network, to powerful cloud computing infrastructures.
Configurability. While offering a standard interface, the implementation and behavior of the Radon runtime can be configured to adapt to various requirements. For instance, two runtime implementations may adopt different compilation techniques for atoms. Ahead-of-Time (AOT) compilation may lead to better and more predictable performance, while Just-in-Time (JIT) compilation may reduce the time to instantiate and start an atom. Likewise, runtime implementations may adopt different communication protocols depending on the underlying network infrastructure, covering the different requirements of the edge-to-cloud continuum deployments.
Figure 1: The Radon model deploys atom instances across a heterogeneous network of hosts (e.g., cloud servers, edge servers, edge, mobile and IoT sensors). Developers write atom definitions and configurations, and the Radon runtime manages the execution of atom instances and their interaction through a runtime interface, with libraries extending core interface functions.
III The Radon Model
Figure 1 shows the high-level architecture of an application developed in Radon. The Radon programming model promotes modularity building on the core concept of atoms. Developers implement their application writing atom definitions, which serve as blueprints and define the behavior of atoms. An atom configuration is then associated with each definition, determining scheduling and deployment choices. By combining an atom definition with its configuration, Radon creates and deploys atom instances across heterogeneous edge-to-cloud environments. Although these instances share the same underlying code (i.e., the definition), each one runs independently in its own memory space and executes concurrently with the others.
Atom instances leverage the services offered by the Radon runtime, which, in analogy with an operating system, plays the role of the kernel that exposes a runtime interface to atoms [^2] To enhance readability and avoid redundancy, the term atom will be used to refer to the general concept. When necessary, explicit distinctions between atom definition and atom instance will be made.. Through this interface, atoms may access a set of runtime functions to control atom execution, communicate with each other (naming and messaging), access storage, set recovery policies, and access additional OS-level features.
Atoms may also internally rely on libraries that extend and enrich the runtime interface. Libraries can be used to abstract features implemented on top of the runtime, which can then be reused across multiple atoms. For instance, a library may implement causally-ordered delivery of messages on top of the FIFO-ordered messaging primitive exposed by the runtime.
In the remainder, we detail the most important functionalities offered by the runtime, such as atom execution (Section III-A), naming and messaging (Section III-B), storage (Section III-C), and recovery (Section III-D). We omit the discussion on other functionalities, such as querying the system’s real-time clock or generating secure random numbers, as they don’t significantly contribute to the relevance of the model.
III-A Atom execution
An atom instance is a stateful, single-threaded entity running within the Radon runtime. It communicates with other instances via message passing and may handle dynamic events generated by the external environment, such as incoming HTTP requests. To encode different interaction patterns and architectural styles, we define two kinds of atoms, differing by the timing of their instantiation:
  • Daemon atoms are statically instantiated and have a fixed, predefined name. Their instances are created at the moment their configuration is provided to the runtime, and they remain active throughout the application’s life-cycle (unless they explicitly exit).
  • Reactive atoms are instantiated dynamically in response to events, with their names generated upon instantiation. Unlike daemon atoms, their instances are not created when the configuration is provided to the runtime but are triggered by events.
While daemon atoms act as well-known named services, proactively launched by the application, the number of instances of reactive atoms and how those instances are spawned is determined by their scheduling policy:
  • one: a single atom instance is created when the first event arrives, and it handles all events directed to it;
  • round-robin: a fixed number of instances are created when the first events arrive, and the subsequent events are distributed among them in a round-robin fashion;
  • on-demand: a new instance is created for each event
  • on-demand-expire: a new instance is created only if none is available, with instances being reused until an expiration criteria is met (time or number of events handled).
Fixed instance count policies (one and round-robin) result in a behavior similar to that of threads in a thread-pool. These atoms are instantiated upon request arrival, handle the request, and then return to a ready state for future requests. However, due to the single-threaded nature of atoms, if an instance is already processing a request, it cannot handle new ones until it completes execution. For example, in a resource-constrained edge serverless environment with a maximum of 8 available cores, configuring a reactive atom with a round-robin policy and a limit of 8 instances ensures that at most 8 requests are processed concurrently, preventing resource exhaustion. Conversely, on-demand and on-demand-expire policies result in a number of instances that scales dynamically according to demand. These policies are well suited for environments with abundant resources, like a serverless infrastructure in a data center, as they enable the system to handle fluctuating workloads without predefined concurrency limits.
The atom configuration defines the nature (daemon vs. reactive) and scheduling policy of the atom, together with the set of hosts where it must be instantiated. It follows that the entire deployment of an application is determined by the set of configurations of its atoms. At the same time, programmers may change the architectural style and interaction pattern of their application by changing the configurations of its atoms.
The current implementation of Radon includes a deployment component that uses atom configurations to bootstrap the runtime and instantiate atoms on a set of known hosts. It also includes a tagging mechanism to whitelist or blacklist hosts, this way an atom can be instantiated only on the hosts that provide the required capabilities.
III-B Naming and messaging
The Radon runtime allows atoms to interact through a message-based communication layer, and it also provides a naming service, allowing atom instances to exchange messages using their names. From the atoms point of view, there is no difference between sending a message to an instance on the same host or to one that runs remotely. On the other hand, the implementation of the runtime may exploit placement information to optimize communication between co-located instances, hiding these implementation details from application developers. As an example, the current Radon implementation bypasses the network stack for local communication, leveraging OS-level memory sharing instead.
To support the most common interaction patterns of modern distributed applications, the Radon naming system offers explicit names, which directly identify atoms, and also aliases that can be used for implicit messaging patterns such as pub-sub. Atoms use these names and aliases without caring about how they are translated to physical addresses, which is an implementation detail hidden behind the Radon runtime interface. In particular, atoms query the naming system to discover instances based on their name, part of their name, or an alias. The current Radon implementation offers regex-based queries to perform such an advanced form of resolution.
Regarding the actual messaging, the Radon runtime exposes both one-to-one and one-to-many messaging facilities. Moreover, messages can be sent either without ordering guarantees or with FIFO ordering guarantees. Send operations are non-blocking: if the destination is a valid atom, the message will be queued and sent as soon as possible. Instead, receiving a message is a synchronous operation: atoms block their execution, waiting for messages.
III-C Storage
Operating systems abstract disks and, in general, the physical storage of nodes using file systems. Radon adopts a similar approach but offers a higher-level interface in the form of a per-node key-value store: atoms can save data to the node’s persistent storage using get and set operations on this store.
The choice of such a simple but effective abstraction aligns with the objectives of Radon: (i) It promotes flexibility, as it can easily be used as a building component to implement more complex data management solutions  [^3] https://tikv.org/. (ii) It grants portability, as it can be easily adapted to different architectures and hardware platforms. Regarding this last aspect, the interface we propose can be implemented on a server, but it can also be deployed using a browser’s local storage or a Database-as-a-Service, adapting to various compute continuum scenarios.
III-D Recovery
Every application may fail due to unexpected circumstances, especially in dynamic environments. The Radon runtime provides mechanisms to handle atom faults and possibly perform recovery operations. In particular, in the atom configuration, it is possible to specify a recovery policy that determines how the runtime behaves in the event of errors occurring at runtime:
  • the None policy (default) signals an exception and stops the atom, reporting the error to the runtime; the rest of the atoms will keep running as if no fault had happened;
  • the escalate policy stops the runtime completely along with all other atom instances running on the same node;
  • the restart policy re-instantiates the atom with the same name;
  • the recover policy executes recovery operations specified by the developer before restarting the instance.
Applications that cannot recover if an important atom fails can use the escalate policy to prevent illegal application states, while applications with recoverable faults can use the restart and recover policies to perform corrective operations (if needed) and continue with the execution.
IV Evaluation
We evaluate our proposal by implementing a concrete application: a distributed and replicated key-value store. Despite being a proof of concept, the application covers all the key features of Radon: nodes use the storage interface to persist data, the communication primitives to redirect requests, and the naming service to enable new nodes to join the network. We use the application to assess both the generality and ease of use of the Radon programming model and to conduct a preliminary assessment of performance.
IV-A A distributed and replicated key-value store in Radon
The implementation of the our key-value store is inspired by Anna [7] and its design focuses on scalability and high availability. The system is organized in three atom definitions: KVNodeD, CoordinatorD, and KVFrontend.
KVNodeD sits at the core of the application. It controls storage access and provides a put-get message-based API for key-value pairs. Multiple KVNodeD instances are organized in a ring topology using consistent hashing. Each instance is responsible for one or more key partitions and replicates data across the ring based on a configurable replication factor $N$. Each key-value pair is replicated across N consecutive elements on the ring.
CoordinatorD manages the ring topology by tracking active KVNodeD instances. A single CoordinatorD instance is deployed per application. It responds to topology queries from other atoms and handles additions of ring elements by notifying neighbors to rebalance key partitions.
KVFrontend is the gateway of the system. It process incoming REST requests for put and get operations in the form of events, and it routes them to the correct KVNodeD instance based on key hashes.
CoordinatorD and KVNodeD are both daemon atoms. The former is instantiated on a single node, while the latter is instantiated multiple times on different hosts. KVFrontend is instead a reactive atom instantiated dynamically to serve incoming requests from clients. It has an on-demand-expire policy with an expiration time of 5s.
At startup, each KVNodeD instance contacts the CoordinatorD instance to register itself and obtain the current ring topology. The CoordinatorD instance updates the topology and notifies all existing KVNodeD instances of the changes, prompting key rebalancing. KVFrontend instances fetch the ring topology from CoordinatorD when instantiated.
Our implementation prioritizes high availability over strong consistency. KVFrontend instances send their request to the KVNodeD responsible for the key to be read or updated. If, due to topology changes, the key is not found at that node, it is responsibility of the receiving KVNodeD to forward the request along the ring, until the first responsible node is found. If the request was a put, the KVNodeD updates its local storage and forwards the request to the following KVNodeD instances until all replicas are updated. If the request was a get, the KVNodeD retrieves the value from its local storage and forwards it back to the requesting KVFrontend.
Atom definitions are written in Rust. As a qualitative measure of code complexity, we implemented shared functions in a common library of $207$ LOC; the KVNodeD, KVFrontend, and CoordinatorD definitions consist of $133$, $88$, and $35$ LOC each, for a total of only $463$ LOC. Despite being a proof of concept far from being a real-world product, these numbers testify that Radon enables developing working distributed applications with low effort.
IV-B Performance assessment
We conduct a performance assessment by deploying multiple nodes of our key-value store and simulating concurrent client requests. We measure performance in terms of throughput and latency. We run the experiments on a Workstation equipped with an AMD Ryzen 9 9950X, 64GB of DDR5 RAM, and a 1TB NVMe PCI Gen3$\times$4 DRAM-less QLC SSD, running Fedora Server 41 (kernel 6.13.5). For consistency, we disable Turbo Boost, limiting the CPU clock to 4.3GHz. We compile all Rust code in release mode with rustc 1.85.0 (stable). The Radon runtime uses Wasmer 5.0.4 with LLVM 18.1.8 to compile the atoms using an AoT compilation strategy.
We simulate a distributed scenario by running three instances of the Radon runtime on different ports (we will refer to these three instances as nodes). Each node hosts the components of the key-value store, specifically 8 KVNodeD instances and the on-demand KVFrontend instances. Additionally, one of the nodes runs a single CoordinatorD. A multi-threaded client submits requests to all three nodes. We emulate multiple clients using asynchronous tasks that send requests sequentially, yielding to the asynchronous runtime while waiting for network operations. Each emulated client randomly selects one of the three nodes and directs all its requests to it. Requests are distributed as follows: 20% put, 40% get (random key), 40% get (recently inserted key). Each client limits its requests to reach a target rate per client, if the rate cannot be reached it sends as many requests as it can.
We measure the throughput in terms of operations per second and the latency of each operation, observed from the client. To contextualize these results, we conduct two additional experiments using the same client workload directed to a dummy Web server that does not store any data, but immediately responds sending back the body of the request. The first baseline experiment, denoted as echo, employs a Web server implemented in vanilla Rust. The second experiment, radon-echo, uses a Web server built in Radon, where an atom is instantiated dynamically with an on-demand-expire policy with a 5s expiration time. We use these measurements as a baseline to isolate the overhead of the runtime itself and the overhead of our experimental setup, to gain a better view of the actual performance of the application we built.
We measure the maximum throughput varying the number of clients (1–128) and the rate of requests from each client (100–10k). Each test runs for 10s. We report the measurements for a replication factor of 2: upon insertion of an item, that item will be stored in 2 KVNodeD instances before the response is sent back to the client. Fig. 2 shows the results we measure: our key-value store can sustain up to about 280k requests/s served across all clients. Further increasing the input rate does not lead to a higher throughput, as requests get queued at the client as it waits for previous responses. In comparison, the baseline radon-echo server reaches about 540k requests/s, while the echo server achieves approximately 700k requests/s. This proves the efficiency of our store: despite the additional forwarding of requests across KVNodeD instances and storage access, the achieved throughput is only a factor of 2 lower than that of a Radon-based server solely echoing HTTP request bodies. Furthermore, the overhead introduced by Radon compared to the vanilla Rust implementation corresponds to a factor of approximately 1.3. This includes the overhead coming from the fact that atoms are compiled to WASM and executed with Wasmer.
To measure latency we run a configuration that does not saturate the system using 16 clients submitting 500 requests/s for 20s. Fig. 3 shows the distribution of the per-request latency observed by the client. Our key-value store exhibits higher latencies compared to both echo and radon-echo. Despite the added overhead of requests forwarding, replication, and storage access, the latency remains within an order of magnitude of the baseline systems. Indeed, Radon manages consistent sub-millisecond latencies on both get and put requests.
Figure 2: Throughput comparison for key-value store, radon-echo, and echo.
Figure 3: Latency distribution across requests for key-value store, radon-echo, and echo.
V Related Work
Different studies have explored the evolution of cloud-centric architectures towards more heterogeneous and decentralized environments, highlighting the potential of these approaches to support latency-sensitive and bandwidth-constrained applications [8, 9]. Menetrey et al. [10] discuss how WASM can serve as a lightweight and portable execution environment for edge-to-cloud computing, a concept that directly reflects Radon’s runtime design. Earlier studies discuss how WASM can be employed outside the browser [11], suggesting its feasibility as a lightweight and portable execution environment for the edge-cloud continuum [10]. The use of WASM as a lightweight virtualization layer was considered in the area of serverless computing: Faasm [12] and Sledge [13] focuses on function performance by leveraging efficient runtime environments, particularly in resource-limited contexts.
Serverless computing is often considered a suitable programming model for the computing continuum. Building on this concept, Serverledge [14] extends the serverless model to edge environments, focusing on minimizing the overhead of function execution. Kumar et al. propose Delta [15] as a high-level scheduler of functions in the continuum, emphasizing the need of ”fluid” computations from edge devices to clouds.
Orthogonal to programming models, several works focus on optimizing resource allocation and deployment strategies in the continuum. Farabegoli et al.[16] extend the pulverization paradigm to support dynamic relocation of components across cloud and edge locations. Similarly, Sedlak et al.[17] provide a middleware that implements an offloading mechanism for computational load redistribution within an edge-fog cluster, taking into account devices’ capabilities.
VI Conclusion and Future Works
In this paper, we presented Radon, a novel programming model and platform designed to facilitate the development of distributed applications across the edge-to-cloud continuum. Radon introduces a flexible abstraction based on atoms, allowing developers to construct complex applications by composing reusable building blocks. It embeds a WASM runtime, enabling components to be executed with limited overhead across the computing continuum. Our prototype implementation shows the feasibility and effectiveness of Radon, as further supported through the development and evaluation of a distributed key-value store.
Looking ahead, there are several directions for future research: (i) We plan to extend the runtime with asynchronous semantics, addressing the current lack of concurrency at the level of individual atom instances. To support more complex interaction patterns among atoms while preserving their single-threaded execution model, we aim to implement a form of implicit cooperation between atoms, similar to interrupt handling in operating systems. (ii) We intend to broaden Radon’s interaction model beyond REST-based request-response patterns. This includes enabling Radon to proactively interact with the external world, not merely as a reaction to incoming events. (iii) We seek to explore methods and techniques to simplify and automate the deployment and dynamic adaptation of distributed applications, particularly in large-scale and heterogeneous environments. (iv) We aim to build a catalog of components on top of Radon to assess the generality of the proposed API. (v) We aim at assessing the performance of the platform in different computing continuum scenarios.
By addressing these challenges, we believe Radon has the potential to become a tool for building next-generation distributed applications, empowering developers with greater flexibility and efficiency across cloud and edge environments.

### v2

@@
 In this approach, applications are developed as a composition of loosely-coupled services.
-In fact, managed environments offer application-agnostic platforms with advanced deployment and execution models, such as serverless functions [2], managed databases [3], or data analytics platforms [4, 5].
+In fact, managed environments offer application-agnostic platforms with advanced deployment and execution models, such as serverless functions [2], managed databases [3], or data analytics platforms [4, 5, 6].
 However, with the expansion of Internet of Things (IoT), smart cities, and industrial automation, new requirements have emerged that challenge conventional cloud solutions.
@@
   3. 3.
-    Providing an initial assessment of the effectiveness and efficiency of Radon through the implementation of a concrete application – a distributed key-value store
+    Providing an initial assessment of the effectiveness and efficiency of Radon through the implementation of a concrete application – a distributed key-value store.
 The remainder of the paper is structured as follows.
-Section II outlines our vision for designing and building Radon.
-Section III describes Radon’s programming and execution model.
-Section IV showcases the application of Radon in implementing a distributed key-value store, and evaluates its performance.
-Finally, Section V reviews related work, and Section VI provides a final overview of the paper, highlighting open research directions.
+Sec.
+II outlines our vision for designing and building Radon.
+Sec.
+III describes Radon’s programming and execution model.
+Sec.
+IV showcases the application of Radon in implementing a distributed key-value store, and evaluates its performance.
+Finally, Sec.
+V reviews related work, and Sec.
+VI provides a final overview of the paper, highlighting open research directions.
 II Vision
@@
 Portability.
-The Radon runtime offers a common substrate for atoms, making them portable across highly heterogeneous environments.
+The Radon runtime offers a common substrate for atoms, making them portable across heterogeneous environments.
 Indeed, the lightweight nature of the runtime makes it suitable to run on diverse hardware architectures, from resource-constrained ones typically found at the edge of the network, to powerful cloud computing infrastructures.
@@
 Likewise, runtime implementations may adopt different communication protocols depending on the underlying network infrastructure, covering the different requirements of the edge-to-cloud continuum deployments.
+III The Radon Model
 Figure 1: The Radon model deploys atom instances across a heterogeneous network of hosts (e.g., cloud servers, edge servers, edge, mobile and IoT sensors).
 Developers write atom definitions and configurations, and the Radon runtime manages the execution of atom instances and their interaction through a runtime interface, with libraries extending core interface functions.
-III The Radon Model
-Figure 1 shows the high-level architecture of an application developed in Radon.
-The Radon programming model promotes modularity building on the core concept of atoms.
+Fig.
+1 shows the high-level architecture of an application developed in Radon.
+The Radon programming model promotes modularity through the core concept of atoms.
 Developers implement their application writing atom definitions, which serve as blueprints and define the behavior of atoms.
@@
 By combining an atom definition with its configuration, Radon creates and deploys atom instances across heterogeneous edge-to-cloud environments.
-Although these instances share the same underlying code (i.e., the definition), each one runs independently in its own memory space and executes concurrently with the others.
-Atom instances leverage the services offered by the Radon runtime, which, in analogy with an operating system, plays the role of the kernel that exposes a runtime interface to atoms [^2] To enhance readability and avoid redundancy, the term atom will be used to refer to the general concept.
-When necessary, explicit distinctions between atom definition and atom instance will be made..
+Although instances share the same underlying code (i.e., the definition), each one runs independently in its own memory space and executes concurrently with the others.
+Atom instances leverage the services offered by the Radon runtime, which, in analogy with an operating system, plays the role of the kernel that exposes a runtime interface to atoms [^2] To enhance readability, we will use the term atom to refer to the general concept, making explicit distinctions between atom definition and atom instance when necessary..
 Through this interface, atoms may access a set of runtime functions to control atom execution, communicate with each other (naming and messaging), access storage, set recovery policies, and access additional OS-level features.
-Atoms may also internally rely on libraries that extend and enrich the runtime interface.
-Libraries can be used to abstract features implemented on top of the runtime, which can then be reused across multiple atoms.
+Atoms may also internally rely on libraries that build on top of the runtime to extend and enrich its interface, and can then be reused across multiple atoms.
 For instance, a library may implement causally-ordered delivery of messages on top of the FIFO-ordered messaging primitive exposed by the runtime.
-In the remainder, we detail the most important functionalities offered by the runtime, such as atom execution (Section III-A), naming and messaging (Section III-B), storage (Section III-C), and recovery (Section III-D).
-We omit the discussion on other functionalities, such as querying the system’s real-time clock or generating secure random numbers, as they don’t significantly contribute to the relevance of the model.
+In the remainder, we detail the most important functionalities offered by the runtime, such as atom execution (Sec.
+III-A), naming and messaging (Sec.
+III-B), storage (Sec.
+III-C), and recovery (Sec.
+III-D).
+We omit the discussion on other functionalities, such as querying the system’s real-time clock or generating secure random numbers, as they do not significantly contribute to the relevance of the model.
 III-A Atom execution
@@
   • round-robin: a fixed number of instances are created when the first events arrive, and the subsequent events are distributed among them in a round-robin fashion;
-  • on-demand: a new instance is created for each event
-  • on-demand-expire: a new instance is created only if none is available, with instances being reused until an expiration criteria is met (time or number of events handled).
+  • on-demand: a new instance is created for each event;
+  • on-demand-expire: a new instance is created only if none is available, with instances being reused until an expiration criterion is met (time or number of events handled).
 Fixed instance count policies (one and round-robin) result in a behavior similar to that of threads in a thread-pool.
@@
 IV-A A distributed and replicated key-value store in Radon
-The implementation of the our key-value store is inspired by Anna [7] and its design focuses on scalability and high availability.
+The implementation of the key-value store is inspired by Anna [8] and its design focuses on scalability and high availability.
 The system is organized in three atom definitions: KVNodeD, CoordinatorD, and KVFrontend.
@@
 It responds to topology queries from other atoms and handles additions of ring elements by notifying neighbors to rebalance key partitions.
-KVFrontend is the gateway of the system. It process incoming REST requests for put and get operations in the form of events, and it routes them to the correct KVNodeD instance based on key hashes.
+KVFrontend is the gateway of the system.
+It processes incoming REST requests for put and get operations in the form of events, and it routes them to the correct KVNodeD instance based on key hashes.
 CoordinatorD and KVNodeD are both daemon atoms.
@@
 If, due to topology changes, the key is not found at that node, it is responsibility of the receiving KVNodeD to forward the request along the ring, until the first responsible node is found.
-If the request was a put, the KVNodeD updates its local storage and forwards the request to the following KVNodeD instances until all replicas are updated.
-If the request was a get, the KVNodeD retrieves the value from its local storage and forwards it back to the requesting KVFrontend.
+If the request is a put, the KVNodeD updates its local storage and forwards the request to the following KVNodeD instances until all replicas are updated.
+If the request is a get, the KVNodeD retrieves the value from its local storage and forwards it back to the requesting KVFrontend.
 Atom definitions are written in Rust.
@@
 IV-B Performance assessment
-We conduct a performance assessment by deploying multiple nodes of our key-value store and simulating concurrent client requests.
-We measure performance in terms of throughput and latency.
+To conduct a performance assessment, we deployed multiple nodes of our key-value store and simulated concurrent client requests, and we measured throughput and latency of requests.
 We run the experiments on a Workstation equipped with an AMD Ryzen 9 9950X, 64GB of DDR5 RAM, and a 1TB NVMe PCI Gen3$\times$4 DRAM-less QLC SSD, running Fedora Server 41 (kernel 6.13.5).
@@
 Additionally, one of the nodes runs a single CoordinatorD.
-A multi-threaded client submits requests to all three nodes.
+A multithreaded client submits requests to all three nodes.
 We emulate multiple clients using asynchronous tasks that send requests sequentially, yielding to the asynchronous runtime while waiting for network operations.
@@
 Furthermore, the overhead introduced by Radon compared to the vanilla Rust implementation corresponds to a factor of approximately 1.3.
-This includes the overhead coming from the fact that atoms are compiled to WASM and executed with Wasmer.
+This includes the overhead of compiling atoms to WASM and executing the with Wasmer.
+Figure 2: Throughput comparison for key-value store, radon-echo, and echo.
 To measure latency we run a configuration that does not saturate the system using 16 clients submitting 500 requests/s for 20s.
@@
 Indeed, Radon manages consistent sub-millisecond latencies on both get and put requests.
-Figure 2: Throughput comparison for key-value store, radon-echo, and echo.
-Figure 3: Latency distribution across requests for key-value store, radon-echo, and echo.
+Figure 3: Latency distribution for key-value store, radon-echo, and echo.
 V Related Work
@@
 Our prototype implementation shows the feasibility and effectiveness of Radon, as further supported through the development and evaluation of a distributed key-value store.
-Looking ahead, there are several directions for future research: (i) We plan to extend the runtime with asynchronous semantics, addressing the current lack of concurrency at the level of individual atom instances.
-To support more complex interaction patterns among atoms while preserving their single-threaded execution model, we aim to implement a form of implicit cooperation between atoms, similar to interrupt handling in operating systems.
+Looking ahead, there are several directions for future research: (i) To address the lack of concurrency at the level of individual atom instances while preserving their single-threaded execution model, we plan to implement a form of asynchronous semantics with implicit cooperation between atoms, similar to interrupt handling in operating systems.
 (ii) We intend to broaden Radon’s interaction model beyond REST-based request-response patterns.
 This includes enabling Radon to proactively interact with the external world, not merely as a reaction to incoming events.
-(iii) We seek to explore methods and techniques to simplify and automate the deployment and dynamic adaptation of distributed applications, particularly in large-scale and heterogeneous environments.
-(iv) We aim to build a catalog of components on top of Radon to assess the generality of the proposed API.
-(v) We aim at assessing the performance of the platform in different computing continuum scenarios.
-By addressing these challenges, we believe Radon has the potential to become a tool for building next-generation distributed applications, empowering developers with greater flexibility and efficiency across cloud and edge environments.
+(iii) We will explore methods and techniques to simplify and automate the deployment and dynamic adaptation of distributed applications, particularly in large-scale and heterogeneous environments.
+(iv) To assess the generality and performance of Radon, we plan to build a catalog of components and by testing their behavior in different computing continuum scenarios.
+We believe Radon has the potential to become a tool for building next-generation distributed applications, empowering developers with greater flexibility and efficiency across cloud and edge environments.
+Acknowledgements
+We acknowledge financial support from the PNRR MUR project PE0000021-E63C22002160007-NEST.